ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: not given Pourkazemi, Mohammad H., University of Shahid Beheshty, Iranian Organization for Educational Evaluation, Iran The impact of teaching maths on the success of the students of economy Math plays a more significant role in economics today than it did in the past. Math is an important means through which one can deal with the economic issues and express them eloquently. Thus, calculus is to be cosidered as an important course to be offered in Economic majors. To express economic issues, it is necessary to implement linear algebra, ordinary differential equations, partial differential equations and difference equations. It this research, we intended to make a survey on the degree of success among the students of economics. The students were classified into two groups. The first group comprised the students with math certificates, whereas the others were graduates of non-math majors. Having defined the criterion variable, we will first try to show that those who are categorized in the former group are more successful than those grouped in the latter. Then, through sampling, and benefitting from correlation and regression coefficients we will show that those students who have obtained higher grades in their calculus courses have shown to be more successful in economics. Those who are in economics have to take courses such as microeconomics, macroeconomics, econometrics, funding and banking operations and resource economics. We will finally conclude that calculus courses will improve the students’ status in economics. They will guarantee the students’ success.